Steering Manufacturing produces snow shovels. The selling price per snow shovel is $29.00. There is no beginning inventory. Costs involved in production are: Direct material $5.00 Direct labor 4.00 Variable manufacturing overhead 4.00 Total variable manufacturing costs per unit $13.00 Fixed manufacturing overhead per year $181,170 In addition, the company has fixed selling and administrative costs of $172,400 per year. During the year, Steering produces 54,900 snow shovels and sells 49,560 snow shovels.

What is cost of goods sold using variable costing?

Variable cost of goods sold $__________
Answer #1

Answer : $ 644,280

Variable cost Per Unit = Direct Material + Direct Labor + Variable Manufacturing Overhead

= 5.00+4.00+4.00 = $ 13.00 Per Unit

Under Varaible costing

Variable Cost of Good sold = units Sold * Varaible cost Per unit = 49,560*13 = $644,280 (Answer)